# Characterizing rational homogeneous spaces via $\Csuperscript\C\C^{*}$-actions

### v1

@@ 3. Actions with small criticality and isolated sink/source
Let us start by showing that, if an equalized action has an isolated extremal fixed point, then the space of $1$-cycles modulo numerical equivalence is generated by the numerical classes of some orbits.
@@
Let $(X,L)$ be a smooth polarized pair endowed with an equalized ${\mathbb{C}}^{*}$-action. Then every curve in $X$ is numerically equivalent to a linear combination with rational coefficients of the class of a curve contained in the sink and classes of closures of minimal orbits. In particular, if the action has an isolated extremal fixed point then ${\mathop{\rm N_{1}}\nolimits}(X)$ is generated by classes of closures of minimal orbits.
@@
In this proof we will use the Chow quotient $\mathcal{C}\!X$ of $X$ by the ${\mathbb{C}}^{*}$-action, which is the projective variety defined as the normalization of the closure in $\mathop{\rm Chow}\nolimits(X)$ of the variety parametrizing general ${\mathbb{C}}^{*}$-invariant $1$-cycles in $X$. We refer the interested reader to [23, Section 4] for details. Let us denote its universal family by:
@@
By [17, II.4.19] every curve in $S$ is algebraically equivalent to a linear combination with rational coefficients of $C^{\prime\prime}$ and of the irreducible components of fibers of $p_{|S}$ (in the quoted statement take $X=S$, $Y=B$ and $Z=C^{\prime\prime}$).
Thus any curve in $q(S)$, and in particular $C$, is algebraically, hence numerically, equivalent to a linear combination with rational coefficients of $q_{*}(C^{\prime\prime})$ and of closures of non-trivial orbits. We conclude observing that, by Lemma 2.3, the numerical class of the closure of an orbit is a sum of numerical classes of closures of minimal orbits. ∎
@@ 3.1. Actions of criticality one with an isolated extremal fixed point
Note first that, since the criticality is $1$, we only have two fixed point components, $Y_{0},Y_{1}$, and every $1$-dimensional orbit is minimal. Since moreover the sink is an isolated point, by Proposition 3.2 the Picard number of $X$ is one; in particular every effective divisor is ample. On the other hand, $\nu^{+}(Y_{1})$ must be equal to one, otherwise the Bend-and-Break Lemma applied to closures of $1$-dimensional orbits passing through a point in the source would contradict the fact that the action is equalized, and that every $1$-dimensional orbit is minimal (cf. [25, Proof of Lemma 4.5]).
@@
In the case of criticality two we will consider two different cases, according to the Picard number of $X$.
3.2. Actions of criticality $\mathbf{2}$ and an isolated extremal fixed point (I)
If $X$ has Picard number one, we will relate the action to a special birational transformation of type $(2,1)$. Those are birational transformations $\psi:{\mathbb{P}}^{n}\dashrightarrow Z\subset{\mathbb{P}}^{N}$, where $Z$ is a smooth variety of Picard number one, given by a linear system of degree two, such that $\psi^{-1}$ is given by a linear system of degree one, and the base locus of $\psi$ is smooth and connected.
In [24, Section 5.1] it was explained how to realize special birational transformations of type (2,1) as birational transformations associated with some equalized ${\mathbb{C}}^{*}$-actions. The following Lemma is a partial converse to that construction.
@@ 3.2. Actions of criticality $\mathbf{2}$ and an isolated extremal fixed point (I)
Denote by $Y_{0}\cup Y_{1}\cup Y_{2}$ the fixed point locus of $X_{\infty}$, where $Y_{0}=\{y_{0}\},Y_{2}$ is the source, and $Y_{1}$ is the inner fixed point locus. Note that, by assumption, $X^{{\mathbb{C}}^{*}}=Y_{0}\cup Y_{1}\cup Y_{2}\cup X_{0}$. Since $Y_{0}$ is an isolated point, it follows (as in the proof of Proposition 3.3) that $\nu^{+}(Y_{1})=1$ and, since $X_{0}=B^{-}(Y_{2})$ is a divisor, we have that $\nu^{-}(Y_{2})=1$. Now consider an ample line bundle $L$ on ${\mathcal{X}}$. Its restriction to $X_{\infty}$ is ample, therefore it is a multiple of the ample generator of $\mathop{\rm Pic}\nolimits(X_{\infty})$, and we may claim that $X_{\infty}$ has criticality $2$ with respect to $L_{|X_{\infty}}$. Since $\mu_{L}(Y_{2})<\mu_{L}(X_{0})$ (in fact, $X_{0}$ is the source of the action on $X$, so it is the only fixed point component of maximal weight), it then follows that the criticality of the action on $(X,L)$ is $3$.
Let now $b:{\mathcal{X}}^{\flat}\to{\mathcal{X}}$ be the blowup of ${\mathcal{X}}$ at $y_{0}$, endowed with the B-type ${\mathbb{C}}^{*}$-action inherited from $X$, whose sink is the exceptional divisor $P$; the remaining fixed point locus is isomorphic to $Y_{1}\cup Y_{2}\cup X_{0}$ so, abusing notation, we denote it by $Y_{1}\cup Y_{2}\cup X_{0}\subset{\mathcal{X}}^{\flat}$. Polarizing ${\mathcal{X}}^{\flat}$ with an ample line bundle of the form $L^{\flat}:=mb^{*}(L)-P$, $m\gg 0$, we may assert that the ${\mathbb{C}}^{*}$-action on $({\mathcal{X}}^{\flat},L^{\flat})$ has criticality three, and it satisfies all the assumptions of Corollary 2.8. We may then claim that the birational map $\psi:P\dashrightarrow X_{0}$ induced by the action is the composition of a smooth blowup $s:W:={\mathcal{G}}\!{\mathcal{X}}_{1,2}\to P$ with center isomorphic to $Y_{1}$ and exceptional divisor $E_{1}$, and a smooth blowdown $d:W\to X_{0}$ with center isomorphic to $Y_{2}$ and exceptional divisor $E_{2}$:
@@
The image in $P$ of $E_{2}$ is $B^{+}(Y_{2})\cap P$, which is a hyperplane; more precisely, it is the exceptional divisor of the blowup $X_{0}^{\flat}$ of $X_{0}$ along $y_{0}$. In particular, if $\ell$ is a general line in $P$ and $\widetilde{\ell}$ is its strict transform in $W$, then $E_{2}\cdot\widetilde{\ell}=1$. On the other hand, obviously, $E_{1}\cdot\widetilde{\ell}=0$.
@@
Moreover, if $\alpha$ is equalized, then $\widetilde{\alpha}$ is equalized, as well.
@@
    A Grassmannian of lines in ${\mathbb{P}}^{n}$, ${\rm A}_{n}(2)$.
@@
    The $10$-dimensional spinor variety ${\rm D}_{5}(5)$.
@@
    The $16$-dimensional Cartan variety ${\rm E}_{6}(1)$.
@@
    A general linear complete intersection of codimension $k\leq 2$ of the Grassmannian ${\rm A}_{4}(2)$.
@@
Each of these two cases appears as a smooth linear section of a homogeneous variety –the Spinor ${\rm D}_{5}(5)$, or the Grassmannian ${\rm A}_{4}(2)$– that we denote by $X^{\prime}$. The ${\mathbb{C}}^{*}$-action on $X$ is the restriction of an equalized ${\mathbb{C}}^{*}$-action of criticality two on $X^{\prime}$ with isolated sink. We denote the fixed point components of this action by $Y_{0}^{\prime}=\{y_{0}\}$, $Y^{\prime}_{1}$ and $Y^{\prime}_{2}$.
@@
Given a general hyperplane $H\subset{\mathbb{P}}(\mathop{\rm H}\nolimits^{0}({L^{\prime}})_{1})$, its pullback to ${\mathbb{P}}(N_{Y^{\prime}_{2}|X})$ is given by a nonzero map $s:{\mathcal{O}}_{Y^{\prime}_{2}}\to N_{Y^{\prime}_{2}|X}$. It meets all the quadrics ${\mathcal{L}}_{y}^{-}$, and the rank of the intersection drops precisely on the set of zeros of the composition:
@@
Assume that $X$ has a beautiful family of rational curves ${\mathcal{M}}$; since every curve in $X$ is algebraically equivalent to a ${\mathbb{C}}^{*}$-invariant cycle, it follows that ${\mathcal{M}}$ is necessarily the family of lines in $X$. We will get to a contradiction by looking at the subfamily ${\mathcal{M}}_{y}$ of lines in $X$ passing by a point $y\in Y_{2}$, and the fixed point component ${\mathcal{M}}_{y}^{-}$, which must be smooth, irreducible, and of the expected dimension, if ${\mathcal{M}}$ is beautiful. The above argument tells us that for the points $y$ contained in a linear subspace of codimension $k$ of $Y^{\prime}_{2}$, the variety ${\mathcal{M}}_{y}^{-}={{\mathcal{L}}}_{y}^{-}\cap{\mathbb{P}}(V)$ is singular, a contradiction. ∎
3.3. Actions of criticality $\mathbf{2}$ and an isolated extremal fixed point (II)
@@
Denote by $Y_{0}=\{y_{0}\}$ the sink, by $Y_{2}$ the source, and by $Y_{1}$ the inner fixed point locus of the action, with irreducible components $Y_{1,i}$, $i=1,\dots,k$. The assumptions on $L$ allow us to guarantee that every $1$-dimensional orbit with a limiting point in $Y_{1}$ has $L$-degree one (and so it is minimal), and that there are no orbits linking two different components of $Y_{1}$; in particular, two different $B^{-}(Y_{1,i})$’s can only meet along points of the source $Y_{2}$.
Let us start by assuming that $Y_{2}$ is a divisor in $X$. Since the sink is isolated we know that $\nu^{+}(Y_{1})=1$ (see [25, Proof of Lemma 4.5]), and so $B^{-}(Y_{1})$ is a union of irreducible divisors $E_{i}:=B^{-}(Y_{1,i})$, $i=1,\dots,k$. Moreover, since $Y_{2}$ is a divisor, we have that $\nu^{-}(Y_{2})=1$, and so the union of the $E_{i}$’s is disjoint. Polarizing $X$ with an ample line bundle $L^{\prime}$ assigns weights to the components $Y_{1,i}$, not necessarily equal. The arguments of [25, Section 3] allow us to claim that there is a sequence of morphisms
@@
that contract unions of divisors $E_{i}$, ordered decreasingly by the $L^{\prime}$-weights of the $Y_{1,i}$’s. The final product is a variety $X^{\prime}:=X_{r}$ endowed with an equalized ${\mathbb{C}}^{*}$-action with isolated sink, divisorial source $Y_{2}^{\prime}$ and no inner fixed point components; in particular it has criticality one and we may assert by Proposition 3.3 that $X^{\prime}$ is the projective space ${\mathbb{P}}^{\dim X}$ and $Y_{2}^{\prime}$ is a hyperplane. We conclude that the contraction $\varphi:X\to X^{\prime}$ is a blowup of $X^{\prime}={\mathbb{P}}^{\dim X}$ along a disjoint union of subvarieties of a hyperplane $H\subset{\mathbb{P}}^{\dim X}$.
@@
Assume now that the codimension of the source $Y_{2}$ in $X$ is at least two. Since we are assuming that the Picard number of $X$ is not $1$, by Proposition 3.2 there are closures of orbits $C_{1}$, linking $y_{0}$ and $Y_{1}$, and $C_{2}$, linking $Y_{1}$ and $Y_{2}$, which are not numerically proportional to the closure $C_{g}$ of the general orbit. Denote by $V^{1}$ and $V^{2}$ the normalization of the components of $\mathop{\rm RatCurves}\nolimits(X)$ containing the elements parametrizing $C_{1}$ and $C_{2}$. Since $L\cdot C_{i}=1$ the families are unsplit. Using formula (6) we compute
$-K_{X}\cdot C_{1}-K_{X}\cdot C_{2}=-K_{X}\cdot C_{g}=2\dim X-\dim Y_{2}\geq\dim X+2.$

### v2

@@ 3. Actions with small criticality and isolated sink/source
Let us start by showing that, if an equalized ${\mathbb{C}}^{*}$-action has an isolated extremal fixed point, then the space of $1$-cycles modulo numerical equivalence is generated by the numerical classes of some orbits.
@@
Let $(X,L)$ be a smooth polarized pair endowed with an equalized ${\mathbb{C}}^{*}$-action. Then every curve in $X$ is numerically equivalent to a linear combination with rational coefficients of the class of a curve contained in the sink and classes of closures of minimal orbits. In particular, if the action has an isolated extremal fixed point, then ${\mathop{\rm N_{1}}\nolimits}(X)$ is generated by classes of closures of minimal orbits.
@@
In this proof we will use the Chow quotient $\mathcal{C}\!X$ of $X$ by the ${\mathbb{C}}^{*}$-action, which is the projective variety defined as the normalization of the closure in $\mathop{\rm Chow}\nolimits(X)$ of the variety parametrizing general ${\mathbb{C}}^{*}$-invariant $1$-cycles in $X$. We refer the interested reader to [24, Section 4] for details. Let us denote its universal family by:
@@
By [17, II.4.19], every curve in $S$ is algebraically equivalent to a linear combination with rational coefficients of $C^{\prime\prime}$, and of the irreducible components of fibers of $p_{|S}$ (in the quoted statement take $X=S$, $Y=B$ and $Z=C^{\prime\prime}$).
Thus any curve in $q(S)$, and in particular $C$, is algebraically, hence numerically, equivalent to a linear combination with rational coefficients of $q_{*}(C^{\prime\prime})$ and of closures of $1$-dimensional orbits. We conclude observing that, by Corollary 2.4, the numerical class of the closure of an orbit is a sum of numerical classes of closures of minimal orbits. ∎
@@ 3.1. Actions of criticality one with an isolated extremal fixed point
Note first that, since the criticality is $1$, we only have two fixed point components, $Y_{0},Y_{1}$, and every $1$-dimensional orbit is minimal. Since moreover the sink is an isolated point, by Proposition 3.2 the Picard number of $X$ is one; in particular every effective divisor is ample. On the other hand, $\nu^{+}(Y_{1})$ must be equal to one, otherwise the Bend-and-Break Lemma applied to closures of $1$-dimensional orbits passing through a point in the source would contradict the fact that the action is equalized, and that every $1$-dimensional orbit is minimal (cf. [23, Proof of Lemma 4.5]).
@@
Now we will study equalized actions of criticality two with an isolated fixed point; we will consider two different cases, according to the Picard number of $X$.
3.2. Actions of criticality 2 and an isolated extremal fixed point (I)
If $X$ has Picard number one, we will relate the action to a special birational transformation of type $(2,1)$. These are birational transformations $\psi:{\mathbb{P}}^{n}\dashrightarrow Z\subset{\mathbb{P}}^{N}$, where $Z$ is a smooth variety of Picard number one, given by a linear system of degree two, such that $\psi^{-1}$ is given by a linear system of degree one, and the base locus of $\psi$ is smooth and connected.
In [25, Section 5.1] it was explained how to realize special birational transformations of type (2,1) as birational transformations associated with some equalized ${\mathbb{C}}^{*}$-actions. The following Lemma is a partial converse to that construction.
@@ 3.2. Actions of criticality 2 and an isolated extremal fixed point (I)
Denote by $Y_{0}\cup Y_{1}\cup Y_{2}$ the fixed point locus of $X_{\infty}$, where $Y_{0}=\{y_{0}\},Y_{2}$ is the source, and $Y_{1}$ is the inner fixed point locus. Note that, by assumption, ${\mathcal{X}}^{{\mathbb{C}}^{*}}=Y_{0}\cup Y_{1}\cup Y_{2}\cup X_{0}$. Since $Y_{0}$ is an isolated point, it follows (as in the proof of Proposition 3.3) that $\nu^{+}(Y_{1})=1$ and, since $X_{\infty}=B^{+}(Y_{2})$ is a divisor, we have that $\nu^{-}(Y_{2})=1$. Now consider an ample line bundle $L$ on ${\mathcal{X}}$. Its restriction to $X_{\infty}$ is ample, therefore it is a multiple of the ample generator of $\mathop{\rm Pic}\nolimits(X_{\infty})$, and we may claim that $X_{\infty}$ has criticality $2$ with respect to $L_{|X_{\infty}}$. Since $\mu_{L}(Y_{2})<\mu_{L}(X_{0})$ (in fact, $X_{0}$ is the source of the action on ${\mathcal{X}}$, so it is the only fixed point component of maximal weight), it then follows that the criticality of the action on $({\mathcal{X}},L)$ is $3$.
Let now $b:{\mathcal{X}}^{\flat}\to{\mathcal{X}}$ be the blowup of ${\mathcal{X}}$ at $y_{0}$, endowed with the B-type ${\mathbb{C}}^{*}$-action inherited from ${\mathcal{X}}$, whose sink is the exceptional divisor $P$; the remaining fixed point locus is isomorphic to $Y_{1}\cup Y_{2}\cup X_{0}$ so, abusing notation, we denote it by $Y_{1}\cup Y_{2}\cup X_{0}\subset{\mathcal{X}}^{\flat}$. Polarizing ${\mathcal{X}}^{\flat}$ with an ample line bundle of the form $L^{\flat}:=mb^{*}(L)-P$, $m\gg 0$, we may assert that the ${\mathbb{C}}^{*}$-action on $({\mathcal{X}}^{\flat},L^{\flat})$ has criticality three, and it satisfies all the assumptions of Corollary 2.8. We may then claim that the birational map $\psi:P\dashrightarrow X_{0}$ induced by the action is the composition of a smooth blowup $s:W:={\mathcal{G}}\!{\mathcal{X}}_{1,2}\to P$ with center isomorphic to $Y_{1}$ and exceptional divisor $E_{1}$, and a smooth blowdown $d:W\to X_{0}$ with center isomorphic to $Y_{2}$ and exceptional divisor $E_{2}$:
@@
The image in $P$ of $E_{2}$ is $B^{+}(Y_{2})\cap P$, which can be described as the set of tangent directions (at $y_{0}$) of orbits converging to $y_{0}$ inside $X_{\infty}$, that is, as the hyperplane ${\mathbb{P}}(T_{X_{\infty},y_{0}})$ (which is the exceptional divisor of the blowup $X_{\infty}^{\flat}$ of $X_{\infty}$ along $y_{0}$). In particular, if $\ell$ is a general line in $P$ and $\widetilde{\ell}$ is its strict transform in $W$, then $E_{2}\cdot\widetilde{\ell}=1$. On the other hand, obviously, $E_{1}\cdot\widetilde{\ell}=0$.
@@
Moreover, if $\alpha$ is equalized, then we may also require that $\widetilde{\alpha}$ is equalized.
@@
    a Grassmannian of lines in ${\mathbb{P}}^{n}$, ${\rm A}_{n}(2)$;
@@
    the $10$-dimensional spinor variety ${\rm D}_{5}(5)$;
@@
    the $16$-dimensional Cartan variety ${\rm E}_{6}(1)$;
@@
    a general linear complete intersection of codimension $k\leq 2$ of the Grassmannian ${\rm A}_{4}(2)$;
@@
Each of these two cases appears as a smooth linear section of a homogeneous variety – the Spinor variety ${\rm D}_{5}(5)$, or the Grassmannian ${\rm A}_{4}(2)$ – that we denote by $X^{\prime}$. The ${\mathbb{C}}^{*}$-action on $X$ is the restriction of an equalized ${\mathbb{C}}^{*}$-action of criticality two on $X^{\prime}$ with isolated sink. We denote the fixed point components of this action by $Y_{0}^{\prime}=\{y_{0}\}$, $Y^{\prime}_{1}$ and $Y^{\prime}_{2}$.
@@
Given a general hyperplane $H\subset{\mathbb{P}}(\mathop{\rm H}\nolimits^{0}({L^{\prime}})_{1})$, its pullback to ${\mathbb{P}}(N_{Y^{\prime}_{2}|X})$ is given by a non-zero map $s:{\mathcal{O}}_{Y^{\prime}_{2}}\to N_{Y^{\prime}_{2}|X}$. It meets all the quadrics ${\mathcal{L}}_{y}^{-}$, and the rank of the intersection drops precisely on the set of zeros of the composition:
@@
Assume that $X$ has a beautiful family of rational curves ${\mathcal{M}}$; since every curve in $X$ is algebraically equivalent to a ${\mathbb{C}}^{*}$-invariant cycle, it follows that ${\mathcal{M}}$ is necessarily the family of lines in $X$. We will get to a contradiction by looking at the subfamily ${\mathcal{M}}_{y}$ of lines in $X$ passing by a point $y\in Y_{2}$, and the fixed point component ${\mathcal{M}}_{y}^{-}$, which must be smooth, irreducible, and of the expected dimension, if ${\mathcal{M}}$ is beautiful (see Remark 2.9). The above argument tells us that for the points $y$ contained in a linear subspace of codimension $k$ of $Y^{\prime}_{2}$, the variety ${\mathcal{M}}_{y}^{-}={{\mathcal{L}}}_{y}^{-}\cap{\mathbb{P}}(V)$ is singular, a contradiction. ∎
3.3. Actions of criticality 2 and an isolated extremal fixed point (II)
@@
Denote by $Y_{0}=\{y_{0}\}$ the sink, by $Y_{2}$ the source, and by $Y_{1}$ the inner fixed point locus of the action, with irreducible components $Y_{1,i}$, $i=1,\dots,k$. The assumptions on $L$ allow us to guarantee that every $1$-dimensional orbit with a limit point in $Y_{1}$ has $L$-degree one (and so it is minimal), and that there are no closures of orbits linking two different components of $Y_{1}$. In fact, by means of Corollary 2.4, our assumptions imply that $\mu_{L}(Y_{2})-\mu_{L}(Y_{0})=2$, and that given two components $Y,Y^{\prime}$ linked by a connected chain of closures of $1$-dimensional curves $C_{1}\cup\dots\cup C_{k}$ (so that the sink of $C_{1}$ belongs to $Y$, the source of $C_{k}$ belongs to $Y^{\prime}$, and the source of every $C_{i}$ is the sink of $C_{i+1}$), then $\mu_{L}(Y)<\mu_{L}(Y^{\prime})$; then we infer that $\mu_{L}(Y_{1,i})=\mu_{L}(Y_{0})+1$ for every $i$, and that there are no closures of orbits linking two different components of $Y_{1}$. In particular, two different $B^{-}(Y_{1,i})$’s can only meet along points of the source $Y_{2}$.
Let us start by assuming that $Y_{2}$ is a divisor in $X$. Since the sink is isolated, we know that $\nu^{+}(Y_{1})=1$ (see [23, Proof of Lemma 4.5]), and so $B^{-}(Y_{1})$ is a union of irreducible divisors $E_{i}:=B^{-}(Y_{1,i})$, $i=1,\dots,k$. Moreover, since $Y_{2}$ is a divisor, we have that $\nu^{-}(Y_{2})=1$, and so the union of the $E_{i}$’s is disjoint. Polarizing $X$ with an ample line bundle $L^{\prime}$ assigns weights to the components $Y_{1,i}$, not necessarily equal. The arguments of [23, Section 3] allow us to claim that there is a sequence of morphisms
@@
that contract unions of divisors $E_{i}$, ordered decreasingly by the $L^{\prime}$-weights of the $Y_{1,i}$’s. The final product is a variety $X^{\prime}:=X_{r}$ endowed with an equalized ${\mathbb{C}}^{*}$-action with isolated sink, divisorial source $Y_{2}^{\prime}$, and no inner fixed point components. In particular, it has criticality one and we may assert, by Proposition 3.3, that $X^{\prime}$ is the projective space ${\mathbb{P}}^{\dim X}$ and that $Y_{2}^{\prime}$ is a hyperplane. We conclude that the contraction $\varphi:X\to X^{\prime}$ is a blowup of $X^{\prime}={\mathbb{P}}^{\dim X}$ along a disjoint union of subvarieties of a hyperplane $H\subset{\mathbb{P}}^{\dim X}$.
@@
Assume now that the codimension of the source $Y_{2}$ in $X$ is at least two. Since we are assuming that the Picard number of $X$ is not $1$, by Proposition 3.2 there are closures of orbits $C_{1}$, linking $y_{0}$ and $Y_{1}$, and $C_{2}$, linking $Y_{1}$ and $Y_{2}$, which are not numerically proportional to the closure $C_{g}$ of the general orbit. Denote by $V^{1}$ and $V^{2}$ the normalization of the components of $\mathop{\rm RatCurves}\nolimits(X)$ containing the elements parametrizing $C_{1}$ and $C_{2}$. Since $L\cdot C_{i}=1$ the families are unsplit. Using Corollary 2.4 we compute
$\begin{array}[]{r}-K_{X}\cdot C_{1}-K_{X}\cdot C_{2}=\big{(}\mu_{-K_{X}}(Y_{1})-\mu_{-K_{X}}(y_{0})\big{)}-\big{(}\mu_{-K_{X}}(Y_{2})-\mu_{-K_{X}}(Y_{1})\big{)}=\\[3.0pt] =\mu_{-K_{X}}(Y_{2})-\mu_{-K_{X}}(y_{0})=-K_{X}\cdot C_{g}=2\dim X-\dim Y_{2}\geq\dim X+2.\end{array}$